# RadarGaussianDet3D: An Efficient and Effective Gaussian-based 3D Detector with 4D Automotive Radars

### v1

@@ IV-B Comparison with State-of-the-Arts
Results on TJ4DRadSet. Table I reports the detection accuracy and inference speed of different models on TJ4DRadSet [7], including the proposed RadarGaussianDet3D. It achieves accuracy comparable to the state-of-the-art MAFF-Net [12] ($0.3\%~\text{mAP}\text{3D},+0.4\%~\text{mAP}\text{BEV}$), while delivering substantially higher inference speed. This efficiency is maintained even on a lower-performance GPU, leaving room for future integration of more computationally intensive enhancements.
@@
Inference Speed Considerations. Musiat et al. [5] have benchmarked several models (including PointPillars [3]) across GPUs and embedded devices on the VoD [8] dataset, which enables fairer comparison of inference speeds of different models. From Table II, it can be inferred that models with frame rates significantly below that of PointPillars [3] are unlikely to achieve real-time performance (i.e., FPS $10\mathrm{Hz}$) on embedded platforms such as NVIDIA AGX Xavier. Consequently, although MAFF-Net [12] yields the highest accuracy, it is unlikely to meet real-time requirements in practical deployments. In contrast, RadarGaussianDet3D operates even faster than PointPillars [3], suggesting strong potential for real-time deployment on embedded devices in autonomous driving systems.
TABLE III: Ablation Studies of Components on TJ4DRadSet test set
Enhancements 3D AP (%) BEV AP (%)
GS LFA GFA BGL Car Ped. Cyc. Tru. mAP Car Ped. Cyc. Tru. mAP
(a) 21.76 25.87 55.53 13.00 29.04 38.62 27.64 60.51 23.97 37.68
(b) 23.05 23.76 57.45 19.30 30.89 36.11 27.63 61.90 28.71 38.59
(c) 22.99 27.94 61.50 16.39 32.20 39.53 30.20 64.84 29.45 41.00
(d) 22.80 27.04 61.83 17.63 32.33 34.89 28.96 67.22 26.81 39.47
(e) 25.77 28.08 64.55 15.61 33.50 39.34 30.57 68.30 26.41 41.15
(f) 26.69 28.18 65.84 19.63 35.08 41.66 30.28 69.62 26.35 41.98
TABLE IV: Ablation Studies of Predicted Gaussian Attributes on TJ4DRadSet test set
@@
IV-C Ablation Study
This section presents ablation studies on TJ4DRadSet [7] to evaluate the effectiveness of the proposed components.
Ablations on PGE and BGL. Table III reports the performance of models obtained by gradually adding modules from the CenterPoint-Pillar [33] baseline (a) to the complete RadarGaussianDet3D (f). Replacing the pillar encoder with Gaussian splatting increases the 3D mAP by $0.1\%$, since each point contributes to multiple BEV grids, yielding denser feature maps. Adding LFA/GFA brings additional gains of $1.3\%$ and $1.5\%~\text{mAP}\text{3D}$, respectively, demonstrating that point-level feature interaction benefits Gaussian attribute prediction. When combined to form the proposed PGE, the performance reaches $33.50\%~\text{mAP}_{\text{3D}}$, indicating that LFA and GFA capture complementary information. Finally, the inclusion of BGL further improves accuracy by $1.6\%$, attributable to its comprehensive measurement of bounding box differences.
Figure 5: Visualization of BEV feature maps from CenterPoint-Pillar [33] (left) and RadarGaussianDet3D without BGL (right) on the TJ4DRadSet test set. The first row shows the reference image, the second row presents the initial BEV feature map after the pillar encoder or PGE, the third row displays the BEV feature map output by the backbone and neck, and the last row shows the final detection results.
Ablations on Predicted Gaussian Attributes in PGE. Table IV presents results for different predicted Gaussian attributes, with all models trained without BGL. Predicting position offsets and opacities leads to performance degradation, confirming that excessive degrees of freedom for Gaussian primitives from sparse radar points are detrimental. Optimal performance is obtained when only scales and rotations are estimated.
@@ V Conclusion
This paper presented RadarGaussianDet3D, a fast and accurate 4D radar-based 3D object detector. To mitigate the inherent sparsity of radar point clouds, the proposed Point Gaussian Encoder (PGE) replaces the conventional pillar encoder by predicting Gaussian primitives from aggregated point features and applying 3D Gaussian Splatting to generate dense BEV feature maps. In addition, the Box Gaussian Loss (BGL) comprehensively measures bounding box differences by accounting for correlations among attributes through box-to-Gaussian conversion. Extensive experiments demonstrated that RadarGaussianDet3D achieves state-of-the-art accuracy while delivering substantially lower inference latency, highlighting its suitability for real-time autonomous driving applications.
By unifying radar points and 3D bounding boxes under Gaussian-based representations, this work contributes a new perspective to 4D radar-based 3D object detection. Future research will extend this paradigm to multi-modal fusion and temporal modeling for enhanced perception in complex driving scenarios.

### v2

@@ IV-B Comparison with State-of-the-Arts
Results on TJ4DRadSet. Table I reports the detection accuracy and inference speed of different models on TJ4DRadSet [36], including the proposed RadarGaussianDet3D. It achieves accuracy comparable to the state-of-the-art MAFF-Net [3] ($0.3\%~\text{mAP}_{\text{3D}},+0.4\%~\text{mAP}_{\text{BEV}}$), while delivering substantially higher inference speed. This efficiency is maintained even on a lower-performance GPU, leaving room for future integration of more computationally intensive enhancements.
@@
Inference Speed Considerations. Musiat et al. [20] have benchmarked several models (including PointPillars [14]) across GPUs and embedded devices on the VoD [21] dataset, which enables fairer comparison of inference speeds of different models. From Table II, it can be inferred that models with frame rates significantly below that of PointPillars [14] are unlikely to achieve real-time performance (i.e., FPS $10\mathrm{Hz}$) on embedded platforms such as NVIDIA AGX Xavier. Consequently, although MAFF-Net [3] yields the highest accuracy, it is unlikely to meet real-time requirements in practical deployments. In contrast, RadarGaussianDet3D operates even faster than PointPillars [14], suggesting strong potential for real-time deployment on embedded devices in autonomous driving systems. The speed gain of PGE primarily stems from avoiding point padding and redundant computation on empty points in pillar-based encoders, together with the efficiency of Gaussian splatting.
IV-C Ablation Study
This section presents ablation studies to evaluate the effectiveness of the proposed components. Results on TJ4DRadSet [36] dataset are reported by default.
Ablations on PGE and BGL. Table III reports the performance of models obtained by gradually adding modules from the CenterPoint-Pillar [32] baseline (a) to the complete RadarGaussianDet3D (g). Replacing the pillar encoder with Gaussian splatting increases the 3D mAP by $0.1\%$ on TJ4DRadSet, since each point contributes to multiple BEV grids, yielding denser feature maps. Adding LFA/GFA brings additional gains of $1.3\%$ and $1.5\%~\text{mAP}\text{3D}$, respectively, demonstrating that point-level feature interaction benefits Gaussian attribute prediction. When combined to form the proposed PGE, the performance reaches $33.50\%~\text{mAP}_{\text{3D}}$, indicating that LFA and GFA capture complementary information. Finally, the inclusion of BGL further improves accuracy by $1.6\%$, attributable to its comprehensive measurement of bounding box differences.
On VoD [21], a similar overall performance trend is observed, except that detection accuracy decreases from (a) to (b) due to the sparser and more irregular radar point distributions. This observation emphasizes the necessity of the proposed LFA and GFA modules, which effectively recover and further improve performance by aggregating local and global point features.
TABLE III: Ablation Studies of Components on TJ4DRadSet test set and VoD [21] val set
Enhancements TJ4DRadSet 3D AP (%) TJ4DRadSet BEV AP (%) VoD mAP (%)
GS LFA GFA BGL Car Ped. Cyc. Tru. mAP Car Ped. Cyc. Tru. mAP EAA ROI
(a) 23.01 23.32 57.58 19.25 30.79 35.82 26.25 62.65 27.95 38.42 46.84 68.51
(b) 23.05 23.76 57.45 19.30 30.89 36.11 27.63 61.90 28.71 38.59 46.06 67.00
(c) 22.99 27.94 61.50 16.39 32.20 39.53 30.20 64.84 29.45 41.00 48.70 69.12
(d) 22.80 27.04 61.83 17.63 32.33 34.89 28.96 67.22 26.81 39.47 48.94 69.48
(e) 25.77 28.08 64.55 15.61 33.50 39.34 30.57 68.30 26.41 41.15 50.18 69.52
(f) 25.77 23.47 58.42 18.94 31.65 42.44 26.60 63.04 28.20 40.07 48.30 69.52
(g) 26.69 28.18 65.84 19.63 35.08 41.66 30.28 69.62 26.35 41.98 52.04 70.61
Figure 5: Visualization of BEV feature maps from CenterPoint-Pillar [32] (left) and RadarGaussianDet3D without BGL (right) on the TJ4DRadSet test set. The first row presents initial BEV feature map after the pillar encoder or PGE, the second row displays the BEV feature map output by the backbone and neck, and the last row shows the radar point cloud and final detection results.
Ablations on Predicted Gaussian Attributes in PGE. Table IV presents results for different predicted Gaussian attributes, with all models trained without BGL. Predicting position offsets and opacities leads to performance degradation, confirming that excessive degrees of freedom for Gaussian primitives from sparse radar points are detrimental, likely due to insufficient local context and measurement noise inherent to the radar modality. Optimal performance is obtained when the model focuses on estimating scale and rotation attributes.
TABLE IV: Ablation Studies of Predicted Gaussian Attributes
@@ V Conclusion
This paper presented RadarGaussianDet3D, a fast and accurate 4D radar-based 3D object detector. To mitigate the inherent sparsity of radar point clouds, the proposed Point Gaussian Encoder (PGE) replaces the conventional pillar encoder by predicting Gaussian primitives from aggregated point features and applying 3D Gaussian Splatting to generate dense BEV feature maps. In addition, the Box Gaussian Loss (BGL) comprehensively measures bounding box differences by accounting for correlations among attributes through box-to-Gaussian conversion. Extensive experiments demonstrated that RadarGaussianDet3D achieves high accuracy while delivering substantially lower inference latency, highlighting its suitability for real-time autonomous driving applications.
By unifying radar points and 3D bounding boxes under Gaussian-based representations, this work contributes new perspectives to 4D radar-based 3D object detection. Future research will extend this paradigm to multi-modal fusion and temporal modeling for perception in complex scenarios.